Calculate the pH of the following salts:a) 0.12 M NaNO2 b) 0.10 M CH3NH3Cl c) .45 M NaOCl?

1
Expert's answer
2021-02-03T07:08:03-0500

a)NaNO"_2" is a mixture of HNO"_2" and NaOH

P"_k" a of HNO"_2" =3.15

"pH= 7+1/2[Pka+logC]\\
p^H=7+1/2[3.15+log(0.12)]\\
p^H=8.16"


b)CH3NH3Cl is a salt of strong acid and weak base.

pH=7-1/2[Pkb+logC]

p^H=7-1/2[3.34+log(0.1)]\\

p^H= 5.83


c) NaOCl is a salt of strong base and weak acid

pH= 7+1/2[Pka+logC]\\

p^H=7+1/2[7.53+log(0.45)]\\

p^H=10.59